OpenStudy (anonymous):

***I will give you a medal*** If R is the total resistance for a parallel circuit with two resistors of resistance r1 and r2, then . Find the resistance, r1, if the total resistance R is 20 ohms, and r2 is 75 ohms. Round your answer to the nearest ohm if necessary. 1/R = 1/r1+1/r2

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

1/R = 1/r1 + 1/r2 1/20 = 1/x + 1/75 solve for x

OpenStudy (anonymous):

its going to be a negative 55

jimthompson5910 (jim_thompson5910):

no

OpenStudy (anonymous):

idk then

jimthompson5910 (jim_thompson5910):

the LCD is 150x, multiply each fraction by the LCD to clear out the fractions

OpenStudy (anonymous):

i get 7.5 = 150/r1 + 2

jimthompson5910 (jim_thompson5910):

oh sry, i miscalculated the LCD the LCD is actually 300x 300x times 1/20 = 15x 300x times 1/x = 300 300x times 1/75 = 4x

jimthompson5910 (jim_thompson5910):

so 1/20 = 1/x + 1/75 is the same as 15x = 300 + 4x

jimthompson5910 (jim_thompson5910):

now solve 15x = 300 + 4x for x

OpenStudy (anonymous):

15.7

OpenStudy (anonymous):

and round up .. it will be 16 ?

jimthompson5910 (jim_thompson5910):

15x = 300 + 4x 15x - 4x = 300 11x = 300 x = ???

OpenStudy (anonymous):

27

jimthompson5910 (jim_thompson5910):

good